# VisualSpeech: Enhance Prosody with Visual Context in TTS

### v1

@@ Abstract
Text-to-Speech (TTS) synthesis faces the inherent challenge of producing multiple speech outputs with varying prosody from a single text input. While previous research has addressed this by predicting prosodic information from both text and speech, additional contextual information, such as visual features, remains underutilized. This paper investigates the potential of integrating visual context to enhance prosody prediction. We propose a novel model, VisualSpeech, which incorporates both visual and textual information for improved prosody generation. Empirical results demonstrate that visual features provide valuable prosodic cues beyond the textual input, significantly enhancing the naturalness and accuracy of the synthesized speech. Audio samples are available at https://ariameetgit.github.io/VISUALSPEECH-SAMPLES/.
@@
Text-to-speech Synthesis, Visual Features, Prosody
@@ 1 Introduction
With advancements in deep learning [1, 2, 3], text-to-speech (TTS) [4, 5] models have become capable of generating speech that closely mimics human speech. However, despite the high-quality output, the prosody of the generated speech sometimes becomes inappropriate for the context. For example, the generated speech sometimes fails to generated different degree of happiness of the same sentences in different context. This limitation arises because the same text input can produce multiple speech outputs with varying prosodic patterns. This is a well-known challenge in speech synthesis, known as the one-to-many problem.
Previous research has attempted to mitigate the one-to-many problem by incorporating additional input text, including both textual and speech-related inputs. Commonly used text features encompass linguistic elements (such as part-of-speech, word boundaries, and word position), as well as syntactic and semantic aspects. In addition, some models incorporate an extra module to extract latent features, including pitch, energy, and duration, from reference speech [6, 7, 8], capturing both global and local latent variables. Although these studies have learned various prosody information from different aspects and greatly improved prosody, there is still much work to be done.
Beside the text and speech, in various application scenarios, such as video game characters navigating dungeons or engaging in battles, visual information plays a crucial role in accurately predicting prosody. For instance, the speech patterns of these characters are unlikely to resemble those of read speech due to the dynamic visual contexts. There have so far been limited research on the usefulness of visual information in TTS. For instance, [9] successfully generated speech samples with accurate reverberation effects in specific environments by learning physical environment information from video data. However, no previous studies have explored the impact of visual information on prosody in TTS.
Thus, this paper explores the possibility of improving prosody prediction in TTS by means of visual information. It makes three key contributions. First, it demonstrates that visual cues carry valuable prosodic information. Second, it establishes that this visual information complements existing textual features rather than being redundant. Finally, it reveals that integrating both textual and visual inputs leads to substantial improvements in prosody prediction within TTS models.
@@ 2.1 Overview
This paper first verifies that the visual information is predictive of prosody. Then it proved that the visual-based prosodic information is complementary to that extracted from text. Next, it integrates the visual features in a modern TTS model as shown in Figure 1. It illustrates the overall architecture of the proposed model VisualSpeech, which takes FastSpeech2 [6] as its backbone due to its explicit prosody modeling capabilities and widespread adoption in the field.
As the Figure 1(a) shows, the model mainly consists of four standard FastSpeech2 modules( text encoder, variance adaptor, Mel decoder and pre-trained vocoder and 2 new modules (visual encoder and visual-text fusion).
@@ 2.2 The Visual Encoder
The visual encoder shares a similar structure with the text encoder in Fastspeech2. Visual features $\textbf{F}=\{\textbf{f}_{1},\textbf{f}_{2},\dots,\textbf{f}_{n}\}$. These extracted features F are then passed into the visual encoder to generate the visual hidden sequence $\textbf{V}=\{\textbf{v}_{1},\textbf{v}_{2},\dots,\textbf{v}_{n}\}$.
@@ 2.3 The Visual-Text Fusion
Since the sequence lengths of the text encoder output P and the visual encoder output V are different, a simple addition or concatenation approach is infeasible in this scenario. Simple methods like average pooling can be adopted to compress V into one global visual vector which can be added to the text encoder output P. However, local information might be lost if global visual vector is utilized due to the possibility of information over-compression. Thus, a more dedicated module (i.e. the visual-text fusion module) is proposed to better summarize the visual information and to shed light on how different parts of a video are co-related to the text. The proposed visual-text fusion module is illustrated in Figure 1(b). It is designed to effectively capture and leverage the complex interactions between textual and visual modalities. At its core, this module employs cross-attention that enables bidirectional information flow between the two modalities. This design choice is motivated by the observation that prosodic patterns in speech often correlate with both linguistic content and visual context cues.
The fusion architecture consists of two key components: (1) a cross-attention 1 (as shown in Figure 1(d)), and (2) a feedforward network for feature refinement.
$\delta(\mathbf{V},\mathbf{P})=\text{Softmax}\left(\frac{\mathbf{P}\mathbf{V}^{T}}{\sqrt{\mathbf{d}_{\mathbf{v}}}}\right)\mathbf{V}$ (1)
In the cross-attention layer, text features are used as queries to attend to visual features, enabling the model to learn which visual cues are most relevant to specific textual elements and how these cues contribute to prosodic variation. Using the output of the text encoder $\mathbf{P}$ as the query can ensure that the output length of the visual-text fusion module is the same as the length of the phone embedding sequence, thus the output sequence of the module could be used to predict the pitch, energy, and duration of the phone level. $\mathbf{P}$ and $\mathbf{V}$ undergo a dot product operation to obtain attention scores, which are divided by $d_{v}$ (that is, the dimension of $\mathbf{V}$) and then passed through a Softmax function to get the normalized scores/weights. These normalized attention scores are used to obtain a weighted sum of $\mathbf{V}$.
@@ 3.1 Experimental Setup
To effectively assess the impact of visual information on speech performance, it is essential to have a dataset that offers both diverse prosodic and visual data. However, research exploring the role of visual features in speech synthesis from a prosody perspective is still scarce, and existing datasets are limited in scope. While well-known speech synthesis datasets, such as LJSpeech [10] and LibriTTS [11], provide rich textual data, they lack corresponding visual features. On the other hand, datasets that do include visual information, like TED and Ego4D [12], suffer from limitations in prosodic or visual diversity. To address this gap, this paper utilizes the Condensed Movies Dataset (CMD) [13], an open-sourced, large-scale video dataset in English. CMD consists of over 3,605 films, spanning various genres, countries, and decades, making it a valuable resource for studying the integration of visual features in speech synthesis. These videos capture high-quality scenarios such as conversations at funerals or casual chats on the streets, providing a rich array of prosodic and visual variations.
The original duration of the video clips ranges from 10 to 300 seconds. Each clip was split into sentence segments using Whisper [14] to facilitate model training [15]. A subset of the resulting dataset was randomly selected for the experiments in this study, named CMD2, comprising approximately 33 hours of video and speech [15]. This subset includes 70,599 video clips ranges from 2 to 30 seconds along with corresponding speech and text transcriptions.The resulting duration of video clips ranges from 2 to 30 seconds.
@@
This work utilizes the state-of-the-art vocal extraction model MVSEP [16] for vocal separation and the open-source toolkit Resemble-Enhance [17] for speech denoising and enhancement. Objective evaluations were conducted to assess the impact of vocal extraction, speech denoising, and enhancement on the speech quality. Following the methodology in [18], the evaluation metrics include Short-Time Objective Intelligibility (STOI) [19], Wideband Perceptual Estimation of Speech Quality (PESQ) [20], and Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) [21]. Since clean or reference speech is unavailable, reference-free scores for these metrics were estimated using the method in [22]. A random selection of 1,000 samples from the dataset was used to compute these metrics. The results, presented in Table 1, clearly demonstrate that, compared to the raw dataset, vocal extraction yields relative improvements of 7.7%, 19.4%, and 79.4%, while speech denoising and enhancement provide even more substantial improvements, highlighting the effectiveness of both techniques.
@@
The following filtering criteria are applied to remove singing clips from the dataset. First, speech samples with pitch values exceeding a predefined threshold (i.e., 500 Hz) are discarded. Second, we compute the mean and standard deviation of pitch, energy, and duration for each phoneme. Any speech sample with pitch or duration values that fall outside 2.5 times the standard deviation of these statistics is considered an outlier and is removed. After applying these filters, the resulting dataset is prepared for training the TTS model.
@@
Following common practice in [4], we preprocess all speech into 22.05k sample rate, and 80-dimension of mel-spectrogram are used. We used Montreal Forced Aligner (MFA) to perform forced alignment and extract phoneme duration.
3.2 The Preliminary Study I: Prosody Clues in Visual Features
@@
To verify the hypothesis of visual features in prosody modeling, a preliminary experiment is conducted using a FeedForward Neural Network (FFNN) to predict prosodic features, pitch and energy, solely from visual features.
@@
Given a sequence of visual features $\textbf{V}=\{v_{1},v_{2},...,v_{n}\}$ and a sequence of prosody features $\textbf{P}=\{p_{1},p_{2},...,p_{t}\}$, where $t\gg n$, we average every $n/p$ prosody features to obtain a target prosody feature corresponding to each visual feature. As a result, the final prosody feature sequence $\textbf{Pv}=\{pv_{1},pv_{2},...,pv_{n}\}$ aligns with the visual feature sequence. In this setup, the visual feature sequence V is used as the input to the model, and the final prosody sequence Pv serves as the target. As the duration of each visual feature is fixed, there is no need for duration prediction.
@@
In this experiment, two types of visual features were investigated: one extracted from the Omnivore model [23] [^1] https://github.com/facebookresearch/omnivore and the other from the ResNet50 model [24] [^2] https://v-iashin.github.io/video_features/. The FFNN models consist of two hidden layers, each with 256 dimensions and a dropout rate of 0.5. Each model was trained for 400,000 steps with a batch size of 32. The loss function used is mean squared error (MSE). The Adam optimizer [25] was employed for parameter updates, with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a learning rate of $1\times 10^{-5}$.
@@
Model Pitch $\downarrow$ Energy $\downarrow$
Mean-predictor 5951.77 1276.15
@@
3.3 The Preliminary Study II: Visual Features as a Complement to Textual Information
After confirming that visual information alone can predict prosody features, this section aims to investigate whether visual features complement textual features. To this end, we extend the pitch, energy, and duration (PED) predictor in FastSpeech2 [6] to incorporate both text and visual information. We then compare the performance of the model using both types of features with that of the model using only textual features. This comparison allows us to assess the added value of visual features in enhancing prosody prediction.
The text-only PED predictor, following the methodology outlined in [6], relies solely on textual input for prosody prediction. In contrast, the visual-text PED predictor integrates both text and visual features to predict prosody. The prosody features—pitch, energy, and duration (P/E/D)—are extracted from the ground-truth (original) audio and serve as the target for both predictors. The visual encoder shares the same architecture as the text encoder. The cross-attention module in the visual-text fusion model has a hidden size of 256, with two attention heads and a dropout rate of 0.2. The training setup follows that of [6], with the exception of a batch size of 32 and a total of 400,000 training steps. The Adam optimizer [25] is used for parameter updates, with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a learning rate scheduler, as specified in [6].
@@ 3.3 The Preliminary Study II: Visual Features as a Complement to Textual Information
Model Pitch $\downarrow$ Energy $\downarrow$ Duration $\downarrow$
@@ 3.4 TTS Experiments
Previous results demonstrate that the combination of visual and textual information outperforms text-only prosody prediction, including pitch and duration. To verify whether these improvement in prosody feature prediction will lead to better speech generation, visual features are integrated into the FastSpeech2 [6] model  [^3] We implemented experiments based on https://github.com/ming024/FastSpeech2, incorporating a visual encoder and visual-text fusion, to synthesize speech. Two VisualSpeech models were trained, one with Omniviore and another with ResNet50 visual features. Each model is trained for 400,000 steps with a batch size of 32. The Adam optimizer is used for parameter updates [25], with hyper-parameters $\beta_{1}=0.9$, $\beta_{2}=0.98$, and the same learning rate scheduler as [6].
@@
Pitch $\downarrow$ Energy $\downarrow$ Duration $\downarrow$
@@
Table 5: The Mel-cepstral Distance (MCD) (in DB) and Log F0 RMSE Loss of Three Models]
@@
Table 4, the proposed model significantly outperform the baseline Fastspeech2 model on all three metrics. More specifically, the proposed model achieves about 33%, 5%, and 49% relative improvement over the baseline model in pitch, energy, and duration prediction, respectively. We find that the improvement in the energy prediction is relatively less significant, which may be attributed to energy being less predictable and trivial. Likewise, FastPitch [26] finds removing energy prediction while maintaining pitch and duration prediction could generate high-quality speech.
Table 5 shows that the proposed model outperforms the baseline model on Mel-Cepstral Distortion (MCD) by a large margin, while the addition of visual features leads to a small improvement on Log F0.

### v2

# VisualSpeech: Enhancing Prosody Modeling in TTS Using Video

@@ Abstract
Text-to-Speech (TTS) synthesis faces the inherent challenge of producing multiple speech outputs with varying prosody given a single text input. While previous research has addressed this by predicting prosodic information from both text and speech, additional contextual information, such as video, remains under-utilized despite being available in many applications. This paper investigates the potential of integrating visual context to enhance prosody prediction. We propose a novel model, VisualSpeech, which incorporates visual and textual information for improving prosody generation in TTS. Empirical results indicate that incorporating visual features improves prosodic modeling, enhancing the expressiveness of the synthesized speech. Audio samples are available at https://ariameetgit.github.io/VISUALSPEECH-SAMPLES/.
@@
Text-to-speech Synthesis, Video, Visual Features, Prosody
@@ 1 Introduction
With advancements in deep learning [1, 2, 3], text-to-speech (TTS) [4, 5] models have become capable of generating speech that closely mimics human speech. Despite the high-quality output, the prosody of the generated speech sometimes becomes inappropriate for the context. For example, most TTS approaches would fail to convey a different degree of happiness within the same sentence given a different context. This limitation arises because the same text input can produce multiple speech outputs with varying prosodic patterns. This is a well-known challenge in speech synthesis, known as the one-to-many mapping problem [6].
Previous research has attempted to mitigate the one-to-many problem by incorporating additional inputs, including both textual and speech-related features. Commonly used text features encompass linguistic elements (e.g. part-of-speech, word boundaries, and word position) as well as syntactic structures and semantic meanings. In addition, some models incorporate an extra module to extract latent features, including pitch, energy, and duration, from reference speech [7, 8, 9], capturing both global and local latent factors. Although these studies have enhanced prosodic modelling in TTS, there is still much work to be done.
In addition to text and speech, many common application scenarios, such as video game characters navigating dungeons or engaging in battles, feature visual information, which could play a crucial role in accurately predicting prosody. For instance, the speech patterns of these video game characters are unlikely to resemble those of read or acted speech due to the dynamic nature of visual contexts. There have so far been limited research on the usefulness of visual information in TTS. For instance, [10] successfully generated speech samples with accurate reverberation effects in specific environments by learning physical environment information from video data. However, no previous studies have explored the impact of visual information on prosody in TTS.
Thus, this paper explores the possibility of improving prosody prediction in TTS with visual information (video). It makes three key contributions. First, it demonstrates visual cues carry valuable prosodic information. Second, this visual information complements existing textual features. Finally, it reveals that the integration of both textual and visual inputs leads to substantial improvements in prosody prediction in TTS models.
@@ 2.1 Overview
This paper first verifies that visual information is predictive of prosody. It then demonstrates that the prosodic information derived from visual features complements that extracted from text. Lastly, the visual features are integrated into a modern TTS model, as shown in Figure 1. The architecture of the proposed model is based on FastSpeech2 [7] due to its explicit prosody modeling and widespread adoption in the field.
As shown in Figure 1(a), the VisualSpeech model mainly consists of four standard FastSpeech2 modules (text encoder, variance adaptor, Mel-spectrogram decoder, and pre-trained vocoder) and two new modules (visual encoder and visual-text fusion). Please refer to [7] for details on the former modules.
@@ 2.2 The Visual Encoder
The visual encoder shares a similar structure with the text encoder in FastSpeech2. Visual features, which refer to the characteristics extracted from video frames (such as facial expressions, body gestures, speaking environment, and other relevant visual cues), are represented as $\textbf{F}=\{\textbf{f}_{1},\textbf{f}_{2},\ldots,\textbf{f}_{n}\}$. These extracted features F are then passed into the visual encoder to generate the visual hidden sequence $\textbf{V}=\{\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{n}\}$.
@@ 2.3 The Visual-Text Fusion
Since the sequence lengths of the text encoder output T and the visual encoder output V are different, a simple addition or concatenation approach is infeasible in this case. Simple methods like average pooling can be adopted to compress V into one global visual vector which can be added to the text encoder output T. However, local information might be lost if global visual vector is utilized due to the possibility of information over-compression. Thus, a more dedicated module (i.e. the visual-text fusion module) is proposed to better summarize the visual information and to shed light on how different parts of a video are correlated to the text. The proposed visual-text fusion module is illustrated in Figure 1(b). It is designed to capture and leverage the complex interactions between textual and visual modalities. At its core, this module employs cross-attention that enables bidirectional information flow between the two modalities. This design choice is motivated by the observation that prosodic patterns often correlate with both linguistic content and visual context cues. The fusion architecture consists of two key components: (1) a cross-attention
$\delta(\mathbf{V},\mathbf{T})=\text{Softmax}\left(\frac{\mathbf{T}\mathbf{V}^{T}}{\sqrt{{d}_{\mathbf{V}}}}\right)\mathbf{V}$ (1)
which is illustrated in Figure 1(d), and (2) a feedforward network for feature refinement. In the cross-attention layer, text features are used as queries to attend to visual features, enabling the model to learn which visual cues are most relevant to specific textual elements and how these cues contribute to prosodic variation. Using the output of the text encoder $\mathbf{T}$ as the query can ensure that the output length of the visual-text fusion module is the same as the length of the phone embedding sequence, thus the output sequence of the module could be used to predict the pitch, energy, and duration at the phone level. $\mathbf{T}$ and $\mathbf{V}$ undergo a matrix-matrix product operation to obtain attention scores, which are divided by the square root of $d_{\mathbf{V}}$ (i.e. dimensionality of $\mathbf{V}$) and then passed through a Softmax function to get the normalized scores/weights. These normalized attention scores are used to obtain a weighted sum of $\mathbf{V}$.
@@ 3.1 Experimental Setup
To assess the impact of visual information on speech generation performance, a dataset containing both diverse prosodic and visual data is essential. However, research exploring the role of visual features in speech synthesis from a prosody perspective is still scarce, and existing datasets are limited in scope. While well-known speech synthesis datasets, such as LJSpeech [11] and LibriTTS [12], provide rich textual data, they lack corresponding visual features. On the other hand, datasets that do include visual information, like TED and Ego4D [13], suffer from limitations in prosodic or visual diversity. To address this gap, this paper utilizes the Condensed Movies Dataset (CMD) [14], an open-sourced, large-scale video dataset in English. CMD consists of short video clips extracted from over 3,605 films, spanning various genres, countries, and decades, making it a valuable resource for studying the integration of visual features in speech synthesis. These videos capture high-quality scenarios such as conversations at funerals and weddings, providing a rich array of prosodic and visual variations.
The original duration of the video clips ranges from 10 to 300 seconds. Each clip was split into sentence segments by means of Whisper [15] to facilitate model training [16]. A subset of the resulting dataset was randomly selected for the experiments in this study, named CMD2, comprising approximately 33 hours of video and speech [16]. This subset includes 70,599 video clips ranging from 2 to 30 seconds along with corresponding speech and text transcriptions.
@@
This work utilizes MVSEP [17] for vocal separation and Resemble-Enhance [18] for speech denoising and enhancement. Objective evaluations were conducted to assess the impact of vocal extraction, speech denoising, and enhancement on the speech quality. Following the methodology in [19], the evaluation metrics include Short-Time Objective Intelligibility (STOI) [20], Wideband Perceptual Evaluation of Speech Quality (PESQ) [21], and Scale-Invariant Signal-to-Distortion Ratio  (SI-SDR) [22]. As reference (clean) speech is unavailable, reference-free scores for these metrics were estimated using the method in [23]. A random selection of 1,000 samples from the dataset was used to compute these metrics. The results, presented in Table 1, clearly demonstrate that, compared to the raw dataset, vocal extraction yields relative improvements of 7.7%, 19.4%, and 79.4%, while speech denoising and enhancement provide even more substantial improvements, highlighting the effectiveness of both techniques.
@@
The following filtering criteria are applied to remove singing clips from the dataset. First, speech samples with pitch values exceeding a predefined threshold (i.e., 500 Hz) are discarded. Second, we compute the mean and standard deviation of pitch, energy, and duration for each phoneme. Any speech sample with pitch or duration values that fall outside 2.5 times the standard deviation of these statistics is considered an outlier and is removed. After applying these filters, the resulting dataset(44,665 training, 2,000 validation, and 200 test samples) is prepared for training the TTS model.
@@
Following the common practice in [4], we resampled all speech into 22.05k sample rate, and converted it into 80-dimensional Mel-spectrograms. We used the Montreal Forced Aligner (MFA) to perform forced alignment and extract phoneme duration.
3.2 Preliminary Study I: Prosody Clues in Visual Features
@@
To verify the hypothesis of visual features in prosody modeling, a preliminary experiment is conducted using a FeedForward Neural Network (FFNN) to predict prosodic features (pitch and energy), solely from visual features.
@@
Given a sequence of visual features $\mathbf{V}=\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}\}$ and a sequence of prosodic features $\mathbf{P}=\{\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{t}\}$, where $t\gg n$, we average every $t/n$ prosody features to obtain a target prosody feature corresponding to each visual feature. As a result, the final prosody feature sequence $\mathbf{P}_{\mathbf{V}}=\{\mathbf{p}_{\mathbf{v}_{1}},\mathbf{p}_{\mathbf{v}_{2}},\ldots,\mathbf{p}_{\mathbf{v}_{n}}\}$ aligns with the visual feature sequence. In this setup, the visual feature sequence $\mathbf{V}$ is used as the input to the model, and the prosody sequence $\mathbf{P}_{\mathbf{V}}$ serves as the target.
In this experiment, two types of visual features were investigated: one extracted from the Omnivore model [24] [^1] https://github.com/facebookresearch/omnivore and the other from the ResNet50 model [25]. [^2] https://github.com/v-iashin/video_features
The results in the accepted paper contain a typographical error.
@@
The FFNN models consist of two hidden layers, each with 256 dimensions and a dropout rate of 0.5. Each model was trained for 400,000 steps with a batch size of 32. The loss function used is mean squared error (MSE). The Adam optimizer [26] was employed for parameter updates, with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a learning rate of $1\times 10^{-5}$.
@@
Models Pitch $\downarrow$ Energy $\downarrow$
Mean-predictor 2.16 3.31
@@
3.3 Preliminary Study II: Visual Features as a Complement to Textual Information
Having confirmed visual information can predict prosody, this section investigates whether visual features complement textual features. To this end, we extend the pitch, energy, and duration (PED) predictor in FastSpeech2 [7] to incorporate both textual and visual information. We then compare performance of the model using both types of features with that of the model using only textual features. This comparison allows us to assess the added value of visual features in enhancing prosody prediction.
The text-only PED predictor, following the methodology outlined in [7], relies solely on textual input for prosody prediction. In contrast, the visual-textual PED predictor integrates both textual and visual features to predict prosody. The prosodic features—pitch, energy, and duration (P/E/D)—are extracted from the ground-truth (original) audio and serve as the target for both predictors. The visual encoder shares the same architecture as the text encoder. The cross-attention module in the visual-text fusion model has a hidden size of 256, with two attention heads and a dropout rate of 0.2. The training setup follows that of [7], with the exception of a batch size of 32 and a total of 400,000 training steps. The Adam optimizer [26] is used for parameter updates, with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a learning rate scheduler, as specified in [7].
@@ 3.3 Preliminary Study II: Visual Features as a Complement to Textual Information
Models Pitch $\downarrow$ Energy $\downarrow$ Duration $\downarrow$
@@ 3.4 TTS Experiments
Previous results demonstrate that the combination of visual and textual information outperforms text-only prosody prediction, including pitch and duration. To verify whether these improvement in prosody feature prediction will lead to better speech generation, visual features are integrated into the FastSpeech2 [7] model,333 Experiments are implemented based on the open-sourced repository: https://github.com/ming024/FastSpeech2 incorporating a visual encoder and visual-text fusion, to synthesize speech. Two VisualSpeech models were trained, one with Omnivore and another with ResNet50 visual features. Each model is trained for 400,000 steps with a batch size of 32. The Adam optimizer is used for parameter updates [26], with hyper-parameters $\beta_{1}=0.9$, $\beta_{2}=0.98$, and the same learning rate scheduler as [7].
@@
Models Pitch $\downarrow$ Energy $\downarrow$ Duration $\downarrow$
@@
Table 4 shows that the proposed models significantly outperform the baseline FastSpeech2 model on all three metrics. The Omnivore-based model achieves about 33%, 5%, and 49% relative improvement in pitch, energy, and duration prediction, respectively. We find that the improvement in the energy prediction is relatively less significant, which may be attributed to energy being less predictable. FastPitch [27] also found removing energy prediction while maintaining pitch and duration prediction could generate high-quality speech.
Table 5 demonstrates that the proposed model significantly outperforms the baseline model in terms of Mel-Cepstral Distortion (MCD) and UTMOS  [^4] UTMOS https://github.com/sarulab-speech/UTMOS22 results for these models (i.e. 2.91, 3.06 and 3.13, respectively) suggest that these improvements have led to an enhanced perceptual quality. and offers slight improvement in Log F0. Similar trends are observed when using ground truth (GT) PED, suggesting that visual features also influence mel-spectrogram prediction beyond PED. These findings indicate that the current approach extracts some useful prosodic information from video, but there is still significant room for improvement. The remaining gap in MCD suggests that further refinement in the modeling of visual features or a more effective utilization of the available data could enhance their contribution to prosody modeling, extending beyond PED prediction alone.
Table 5: The Mel-cepstral Distance (MCD) (in DB) and Log F0 RMSE Loss of Three Models
@@
FastSpeech2 (+ GT PED) 5.04 0.32
VisualSpeech (Omnivore) (+ GT PED) 4.83 0.30
VisualSpeech (ResNet50) (+ GT PED) 4.82 0.31